you are riding your bike. it takes 28 min to go 8mi. how long will it take you to go 15 mi
Answer by Alan3354(69443)

you are riding your bike. it takes 28 min to go 8mi. how long will it take you to go 15 mi
d = rt
rate = 8/28 miles/minute = 2/7 mi/minute
t = d/r = 15/(2/7) = 105/2
t = 52.5 minutes